A deflection magnetometer is placed with its arms in north-south direction. How and where should a short magnet having M/B baseH = 40 A-m^2/T be placed so that the needle can stay in any position?

Grade:10

1 Answers

Jitender Pal
9 years ago
Sol. According to oscillation magnetometer, T = 2π√I/MB base H ⇒ π/10 = 2π√1.2 * 10^-4/M * 30 * 10^-6 ⇒ (1/20)^2 = 1.2 * 10^-4/M * 30 * 10^-6 ⇒ M = 1.2 *10^-4 *400/30 * 10^-6 = 16 * 10^2 A-m^2 = 1600 A-m^2